OpenStudy (anonymous):

solve the equation In(x+9)-In(x-5)=In(x+7)-In(x-4).

OpenStudy (anonymous):

using the law of logarithms, this can be simplified to \[\ln (x+9/x-5)=\ln (x+7/x-4)\]

OpenStudy (anonymous):

If we exponentiate both sides of the equation by e our new equation will be x+9/x-5=x+7/x-4

OpenStudy (anonymous):

multiply both sides of the equation by x-5 to get x-9=(x+7)(x-5)/(x-4)

OpenStudy (anonymous):

then multiply both sides of the equation by x-4 to get (x-9)(x-4)=(x+7)(x-5)

OpenStudy (anonymous):

do you mean (x+9)?

OpenStudy (anonymous):

woops thanks for that. now we can use the FOIL method to simplify our equation (x+9)(x-4)=x^2+5x-36 (x+7)(x-5)=x^2+2x-35 so our simplified equation is x^2+5x-36=x^2+2x-35

OpenStudy (anonymous):

so there is really no solution for this problem?

OpenStudy (anonymous):

subtract x^2+2x-35 from both sides to get... 3x-1=0 solving for x we get, 3x=1 x=1/3

OpenStudy (anonymous):

okay thanks! That was one of the possibly answers but, the answer key says no solution ?

OpenStudy (anonymous):

I don't see how the answer key says the equation has no solution unless they solved it a different way

OpenStudy (anonymous):

yes, I just think maybe they made a mistake?

OpenStudy (anonymous):

But thank you for helping me! I know how to use your steps for other problems like this! thanks!!!

OpenStudy (anonymous):

Your're very welcome!! xD